OpenStudy (braydonlevi99):

According to Newton's second law of motion, the force acting on an object varies directly with the object's acceleration. Suppose that a force of 27 N results in acceleration of 3 m/s^2. If the object's acceleration becomes 10 m/s^2, what is the force?

OpenStudy (boldjon):

use the first set of numbers to solve for the mass. F=ma m=f/a m= 60/10 = 6kg Use the second set of numbers (the objects mass which we have already solved for remains the same) F=ma a=f/m a = 24N/6kg a = 4 m/s^2

OpenStudy (anonymous):

Well this is more Physics than pure math, but whatevs \[\large \sf F_{net}=m \times a\] \[\large \sf 27~ N = x ~kg \times 3 ~m/s^{2}\]\[\large \sf x~kg = \frac{27~N}{3~m/s^{2}}\]\[\large \sf x~kg=9~kg\] So the object is 9 kg. So if we have \[\large \sf F_{Net}= 9~kg \times 10~m/s^{2}\] what will the Fnet be?

OpenStudy (boldjon):

60 = K*10 K = 6 24 = 6x x = 4m/sec^2

OpenStudy (anonymous):

Just ignore boldjon, I think he has a few screws loose lol

OpenStudy (boldjon):

-_- i know what i'm doing don't listen to this guy

OpenStudy (anonymous):

You're not even doing the right question... lol

OpenStudy (boldjon):

yeh sure